OpenStudy (embracetheweebness98):

Urgent help needed :/ Solve the following system of equations: 4x − 2y − z = −5 x − 3y + 2z = 3 3x + y − 2z = −5

OpenStudy (welshfella):

You can solve these by the method of elimination HinT start by adding equations 2 and 3 This will eliminate z.

OpenStudy (embracetheweebness98):

The middle and lower one?

OpenStudy (welshfella):

yes

OpenStudy (embracetheweebness98):

okay

OpenStudy (embracetheweebness98):

Am I combining them?

OpenStudy (welshfella):

yes add the x terms , add the y terms -2z + 2z = 0 so z disappears. and add 3 and -5.

OpenStudy (embracetheweebness98):

okay

OpenStudy (embracetheweebness98):

4x-2y=-2

OpenStudy (welshfella):

Correct Now we need to eliminate z using 2 different equations You could multiply the first equation by 2 then add it to equation 3.

OpenStudy (embracetheweebness98):

Im confused, didn't we already eliminate z?

OpenStudy (welshfella):

Sorry multiply fisr equation by -2

OpenStudy (welshfella):

Yes but we need another equation in x and y so we have to eliminate z frm another 2 equations.

OpenStudy (welshfella):

so * no 1 by -2:- -8x + 4y + 2z = 10 3x + y - 2z = -5 ADD

OpenStudy (embracetheweebness98):

okay

OpenStudy (embracetheweebness98):

-5x+5y=5

OpenStudy (welshfella):

OK Now we can simplify that one to make our work easier How can we do that?

OpenStudy (embracetheweebness98):

Move a variable to another side?

OpenStudy (welshfella):

No Just divide through by 5 giving us -x + y = 1

OpenStudy (welshfella):

now we have 2 equations in x and y 4x - 2y = -2 -x + y = 1

OpenStudy (welshfella):

Now i realised - we can also divide the first equation by 2 giving 2x - y = 1 -x + y = 1

OpenStudy (welshfella):

* typo 2x - y = -1 -x + y = 1

OpenStudy (embracetheweebness98):

okay

OpenStudy (welshfella):

now add these 2

OpenStudy (embracetheweebness98):

x=0?

OpenStudy (welshfella):

correct

OpenStudy (welshfella):

now plug in x = 0 0 + y = 1 y = 1

OpenStudy (embracetheweebness98):

okay

OpenStudy (welshfella):

Now you can easily find z by plugging in x = 0 and y = 1 into one of the original equations

OpenStudy (embracetheweebness98):

okay

OpenStudy (embracetheweebness98):

-3

OpenStudy (welshfella):

I got 3

OpenStudy (welshfella):

4(0) - 2(1) - z = -5 -2 - z = -5 -z = -3 z = 3

OpenStudy (welshfella):

Its worth checking if the answer is correct by plugging in values of x y and z into another one of the equations.

OpenStudy (welshfella):

try the last one 3(0) + 1 - 2(3) = 1 - 6 = -5 - this checks out

OpenStudy (welshfella):

Hope this has helped gtg right now.

OpenStudy (embracetheweebness98):

Thank you for the help! Yeah, sorry I had to go too, i was running late for a live lesson by the time I finished :/ @welshfella